# Outward Sodium Current, Fine-structure Constant, and Ferroelectric Hysteresis Regimes in the Giant Squid Axon Propagating Action Potential: a Phase Space Approach

### v8

@@ Acknowledgements.
 NKJ wants to thank Prof. J. J. C. Rosenthal making the present work possible by giving him in 1999 the electronic copy of Wa05097a.dat file containing the experimental data collected in Reference [3]. He is greatly indebted to him and to Prof. Francisco Bezanilla for sharing their data. NKJ wants to thank Dr. Bogdan Mihaila for recommending further study of the role and the meaning of temperature independent dimensionless constants $\alpha_{X}$. NKJ wants to thank Dr. Andrew Perlman for asking a question about Hodgkin-Huxley equations close to 55 years ago. NKJ and FC want to thank the Santa Fe Institute for its hospitality during the work on the paper. NKJ and FC want to thank Prof. Geoffrey West, Prof. Benjamin Drukarch and Prof. Avadh Behari Saxena for helpful comments.
@@ Appendix A Clarification of the Definition of Ionic, Membrane, and Capacitive Currents
$J_{ionic}$ = $J_{membrane}$ $J_{capacitive}$
@@ Appendix B Action potential hysteresis loop
The rising edge portion of the polarization current $J_{PM}(V)$ displays two distinct segments. The outgoing current’s quasi-linear part intercepts the current axis with the slope $g_{PM}$approx 56 mS/cm2 and the value of the intercept is $V_{PM}$approx 106.6 mV. Note however that the tangent of the polarization current tends to infinity as the voltage tends to the peak of the action potential, so the quasilinear behavior is only present away from the peak of the action potential. The polarization current at first increases the incoming the total sodiumcurrent and then it decreases it. This implies first an increase and then a steep decrease of sodium’s $M$-channel effective equilibrium potential (Meep). Thus, up to a fraction of a mili-volt from the peak of the action potential, the $M$-channel total ionic current $J_{IM}(V)$ and Meep(V) can be written as follows:
Figure 19: Schematic ferroelectric hysteresis. (a), $(V_{M}-V_{H})$ [mV] vs. $(V_{K}-V_{N})$ [mV]. (b), Schematic ferroelectric hysteresis loop at 4.5 oC, Polarization = $\propto(V_{M}-V_{H})[\propto$ mV] (c), $(V_{K}-V_{N})$ and $(V_{M}-V_{H})$[mV] vs. Temp.[oC]. (d), Schematic ferroelectric hysteresis loop at 19.8 oC, Polarization = $\propto(V_{M}-V_{H})[\propto$ mV] vs. Electric field [$\propto$ mV]. Time direction is given by the arrow $\rightarrow$.
@@ Appendix C Optimum channel density
for up to about 20 ${}^{o}C.$ It is reasonable to assume that the rate constant $\mu_{M}$ has the same dependency on $R_{i}$, $R$, $v$ and $C_{m}$ up to 20 oC as the propagation constant $k$. Using Eq. (33) and the expression Eq. (21d) for $k$, Eq. (32) is written as:
@@ Appendix D Fitting Procedure
Rosenthal-Bezanilla data provides the action potential $V$ at two points along the axon at different temperatures. The action potential data was analyzed at the point furthest away from the stimulus. Rosenthal measured the resistivity at $18.5^{o}$ and determined that his measurement was consistent with the generally used expression for resistivity, also used in this work:
@@
Approximate time rate parameters $\mu_{K}$, $\mu_{M}$ and $\mu_{N}$ and maximum conductance $g_{K}$, $g_{M}$ and $g_{N}$ can be read from Fig. 21, Fig. 8 and Fig. 9; or, the maximum conductance can be determined by Equations (21a), (21b) and (21c) respectively if the corresponding time rate is known (or vice versa). There is no such relation for H-channel currents. Both, the rate $\mu_{H}$ and maximum conductance $g_{H}$ are to be read separately from Fig. 9. Quasilinear segments in phase space are segments of all corresponding ion channels open and as such are amenable to fittings by the usual expressions for ionic currents Eq. 17d, Eq. 17g, and Eq. 17h where the fractions of open channels are fitted by corresponding mAvrami equations describe a completion of a crystallization process from zero to one by the S curve in time. Fittings, by mAvrami equations Eq.(25a), Eq.(25b) and Eq.(25c) are very sensitive to values of time parameters $to_{M}$, $to_{N}$ , $to_{H}$ and the corresponding time rates $\mu_{X}$, and less so to values of $\alpha_{X}$ and $\theta_{X}$. As a result, we have seeded all $\alpha_{X}$ with the value of the fine-structure constant: $\alpha$ = 0.007297352. Also, all three Avrami exponents were seeded, $\theta_{X}$ = 3.78. While the linear segment of recovery sodium current presents itself clearly and intersects the zero current axis below but close to AP peak, the rising edge total incoming sodium doesn’t present a clear quasilinear segment. Inclusion of points beyond the start of negative resistance region in the fit for fraction of open channels results in a quickly diverging Avrami exponent $\theta_{M}$ indicating that incoming sodium current consist of an additional superimposed polarization process in the negative resistance region and up to just before the action potential peak. The fitting region was chosen to be between the inflection point of the action potential (the maximum value of the capacitive current) and the beginning of the negative resistance region. These choices result in the values for sodium’s reversal potential $V_{M}$, time rate $\mu_{M}$ (and $g_{M}$ according to Eq. (21b)) and the time $to_{M}$ for Eq. 17d and (25a). The difference between the experimental current $J_{I}$ and thus obtained sodium current $J_{M}$ yields the polarization current $J_{P}$ = $\Sigma_{i}J_{Pi}$ = $J_{I}$ $J_{M}$.
@@ Appendix E Supplemental Material - Additional Graphs
Figure 23: Currents and linear segments. Note that all currents look almost the same (scale) at the three temperatures (including Fig.  21 and Fig. 8) except the currents $J_{K}$ and $J_{H}$ which increase and decrease relatively as compared with other currents. (a), (c), Capacitive, membrane and ionic currents are displayed with their parsing into mAvrami fits and corresponding polarization currents. The linear slope of the current $J_{K}$ and the corresponding slope of the capacitive current linear segments are the time rate constant $\mu_{K}$ and maximum conductance $g_{K}$. The slope of the linear segment of the ionic current $J_{M}$ is the maximum conductance $g_{M}$ for the sodium M-channel. The corresponding capacitive current linear slope is -$\mu_{M}$ where $\mu_{M}$ is the time rate constant of the M-channel. The polarization currents are discontinuous at the inception of the mAvrami fits of activation currents. (b), (d), The linear slopes of sodium’s and potassium’s capacitive currents are $\mu_{H}$ and -$\mu_{N}$ corresponding to rate constants $\mu_{H}$ and $\mu_{N}$. The linear slopes of $J_{H}$ and $J_{N}$ are $-g_{H}$ and $g_{N}$ where $g_{H}$ and $g_{N}$ are the maximum conductance of the sodium H-channel and potassium N-channel respectively. While slopes of $J_{H}$ and capacitive current $C_{m}\Phi_{H}$ intercept the potential axis at different points, $C_{m}\Phi_{H}$ intercepts the axis at $V\approx V_{M}$.
@@
Figure 26: Rising edge mAvrami fits of fractions of open Mo/M sodium M-channels and fractions of completed $Pio/Pi$ M-channels morphing. Inception and AP peak polarization flips are displayed. Parameters $\alpha_{M}$ and $\alpha_{Pi}$ from Eq.(20) are seeded with the value of the fine-structure constant $\alpha$ = 0.007297352 and parameters $\theta_{M}$, $\theta_{Pi}$ are seeded with the value 3.78. Note: $g_{P1}=g_{P2}$, $V_{P1}=V_{P2}$, and that there is no segment $P3$. (a) The inception polarization segment consists, of two concatenated portions with different time rates. The AP peak polarization segment $P4$ is a single one. Note that polarization time rates $\mu_{P1}$ and $\mu_{P2}$ are close multiples of sodium’s M-channel time rate $\mu_{M}$, but $\mu_{P4}$ is not. (b), (c), (d) Note: $\ln\alpha$ = $\ln(0.007297352...)$ = $-4.920243...$.

### v9

@@ Acknowledgements.
 NKJ wants to thank Prof. J. J. C. Rosenthal making the present work possible by giving him in 1999 the electronic copy of Wa05097a.dat file containing the experimental data collected in Reference [3]. He is greatly indebted to him and to Prof. Francisco Bezanilla for sharing their data. NKJ wants to thank Dr. Bogdan Mihaila for recommending further study of the role and the meaning of temperature independent dimensionless constants $\alpha_{X}$. NKJ wants to thank Dr. Andrew Perlman for asking a question about Hodgkin-Huxley equations close to 55 years ago. NKJ and FC want to thank the Santa Fe Institute for its hospitality during the work on the paper. NKJ and FC want to thank Prof. Geoffrey West, Prof. Benjamin Drukarch and Prof. Avadh Behari Saxena for helpful comments. In particular we are grateful for the encouraging acknowledgement by Prof. Bezanilla’s of the second outward current upon observing the recovery region graph at our poster at the Biophysics Symposium in San Francisco.
@@ Appendix A Clarification of the Definition of Ionic, Membrane, and Capacitive Currents
$J_{ionic}$ = $J_{membrane}$ - $J_{capacitive}$
@@ Appendix B Action potential hysteresis loop
The rising edge portion of the polarization current $J_{PM}(V)$ displays two distinct segments. The outgoing current’s quasi-linear part intercepts the current axis with the slope $g_{PM}$ $\approx$ 56 mS/cm2 and the value of the intercept is $V_{PM}$ $\approx$ 106.6 mV. Note however that the tangent of the polarization current tends to infinity as the voltage tends to the peak of the action potential, so the quasilinear behavior is only present away from the peak of the action potential. The polarization current at first increases the incoming the total sodiumcurrent and then it decreases it. This implies first an increase and then a steep decrease of sodium’s $M$-channel effective equilibrium potential (Meep). Thus, up to a fraction of a mili-volt from the peak of the action potential, the $M$-channel total ionic current $J_{IM}(V)$ and Meep(V) can be written as follows:
Figure 19: Schematic ferroelectric hysteresis. (a), $(V_{M}-V_{H})$ [mV] vs. $(V_{K}-V_{N})$ [mV]. (b), Schematic ferroelectric hysteresis loop at 4.5 oC, Polarization = $\propto(V_{M}-V_{H})[\propto$ mV] (c), $(V_{K}-V_{N})$ and $(V_{M}-V_{H})$[mV] vs. Temp.[oC]. (d), Schematic ferroelectric hysteresis loop at 19.8 oC, Polarization = $\propto(V_{M}-V_{H})[\propto$ mV] vs. Electric field $\propto$ mV]. Time direction is given by the arrow $\rightarrow$.
@@ Appendix C Optimum channel density
for up to about 20 $^{o}C.$ It is reasonable to assume that the rate constant $\mu_{M}$ has the same dependency on $_{i}, RR$, $v$ and $C_{m}$ up to 20 oC as the propagation constant $k$. Using Eq. (33) and the expression Eq. (21d) for $k$, Eq. (32) is written as:
@@ Appendix D Fitting Procedure
Rosenthal-Bezanilla data provides the action potential $V$ at two points along the axon at different temperatures. The action potential data was analyzed at the point furthest away from the stimulus. Rosenthal measured the resistivity at $18.5^{o}$C and determined that his measurement was consistent with the generally used expression for resistivity, also used in this work:
@@
Approximate time rate parameters $\mu_{K}$, $\mu_{M}$ and $\mu_{N}$ and maximum conductance $g_{K}$, $g_{M}$ and $g_{N}$ can be read from Fig. 21, Fig. 8 and Fig. 9; or, the maximum conductance can be determined by Equations (21a), (21b) and (21c) respectively if the corresponding time rate is known (or vice versa). There is no such relation for H-channel currents. Both, the rate $\mu_{H}$ and maximum conductance $g_{H}$ are to be read separately from Fig. 9. Quasilinear segments in phase space are segments of all corresponding ion channels open and as such are amenable to fittings by the usual expressions for ionic currents Eq. 17d, Eq. 17g, and Eq. 17h where the fractions of open channels are fitted by corresponding mAvrami equations describe a completion of a crystallization process from zero to one by the S curve in time. Fittings, by mAvrami equations Eq.(25a), Eq.(25b) and Eq.(25c) are very sensitive to values of time parameters $to_{M}$, $to_{N}$ , $to_{H}$ and the corresponding time rates $\mu_{X}$, and less so to values of $\alpha_{X}$ and $\theta_{X}$. As a result, we have seeded all $\alpha_{X}$ with the value of the fine-structure constant: $\alpha$ = 0.007297352. Also, all three Avrami exponents were seeded, $\theta_{X}$ = 3.78. While the linear segment of recovery sodium current presents itself clearly and intersects the zero current axis below but close to AP peak, the rising edge total incoming sodium doesn’t present a clear quasilinear segment. Inclusion of points beyond the start of negative resistance region in the fit for fraction of open channels results in a quickly diverging Avrami exponent $\theta_{M}$ indicating that incoming sodium current consist of an additional superimposed polarization process in the negative resistance region and up to just before the action potential peak. The fitting region was chosen to be between the inflection point of the action potential (the maximum value of the capacitive current) and the beginning of the negative resistance region. These choices result in the values for sodium’s reversal potential $V_{M}$, time rate $\mu_{M}$ (and $g_{M}$ according to Eq. (21b)) and the time $to_{M}$ for Eq. 17d and (25a). The difference between the experimental current $J_{I}$ and thus obtained sodium current $J_{M}$ yields the polarization current $J_{P}$ = $\Sigma_{i}J_{Pi}$ = $J_{I}$ - $J_{M}$.
@@ Appendix E Supplemental Material - Additional Graphs
Figure 23: Currents and quasi linear segments. Note that all currents look almost the same (scale) at the three temperatures (including Fig.  21 and Fig. 8) except the currents $J_{K}$ and $J_{H}$ which increase and decrease relatively as compared with other currents. (a), (c), Capacitive, membrane and ionic currents are displayed with their parsing into mAvrami fits and corresponding polarization currents. The linear slope of the current $J_{K}$ and the corresponding slope of the capacitive current linear segments are the time rate constant $\mu_{K}$ and maximum conductance $g_{K}$. The slope of the linear segment of the ionic current $J_{M}$ is the maximum conductance $g_{M}$ for the sodium M-channel. The corresponding capacitive current linear slope is -$\mu_{M}$ where $\mu_{M}$ is the time rate constant of the M-channel. The polarization currents are discontinuous at the inception of the mAvrami fits of activation currents. (b), (d), The linear slopes of sodium’s and potassium’s capacitive currents are $\mu_{H}$ and -$\mu_{N}$ corresponding to rate constants $\mu_{H}$ and $\mu_{N}$. The linear slopes of $J_{H}$ and $J_{N}$ are $g_{H}$ and $g_{N}$ where $g_{H}$ and $g_{N}$ are the maximum conductance of the sodium H-channel and potassium N-channel respectively. While slopes of $J_{H}$ and capacitive current $C_{m}\Phi_{H}$ intercept the potential axis at different points, $C_{m}\Phi_{H}$ intercepts the axis at $V\approx V_{M}$.
@@
[…]